OpenStudy (anonymous):

I can't remember how to do this algebra, can someone please help:

OpenStudy (anonymous):

\[\frac{ \sqrt{t} }{ t^{2} }\]

jimthompson5910 (jim_thompson5910):

what do you want to do here?

OpenStudy (anonymous):

well it's for a limit problem, and i need to divide this. . i can't remember how T__T

jimthompson5910 (jim_thompson5910):

what's the full problem?

OpenStudy (anonymous):

\[\lim \frac{ \sqrt{t} +t^{2}}{ 2t-t^{2} } \]

OpenStudy (anonymous):

as t -> infty

OpenStudy (anonymous):

I'm suppose to divide the t^{2} with everything right?

jimthompson5910 (jim_thompson5910):

ok I see what's going on, one sec

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

I would use the idea that \[\large \sqrt{t} = t^{1/2}\]

jimthompson5910 (jim_thompson5910):

to turn \[\large \frac{ \sqrt{t} }{ t^{2} }\] into \[\large \frac{ t^{1/2} }{ t^{2} }\] \[\large t^{1/2-2} \] \[\large t^{1/2-4/2} \] \[\large t^{-3/2} \]

OpenStudy (anonymous):

i see, is the rule to subtract exponents?

OpenStudy (anonymous):

oh wow i remember now thanks so much!

jimthompson5910 (jim_thompson5910):

see where to go from there? if so, then what's the final answer?

OpenStudy (anonymous):

gonna need a minute lol

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

hey jim, the book says -1 , but I got positive one. my answer is: \[\frac{ \sqrt{\frac{ -3 }{ t }}+1 }{ \frac{ 2 }{ t } + 1}\]

OpenStudy (anonymous):

when the denominator is has the variable, then it should be zero right?

OpenStudy (anonymous):

shouldn't it be 1/1?

jimthompson5910 (jim_thompson5910):

\(\large 2t-t^{2}\) in the denominator should turn into \(\large \frac{ 2 }{ t } - 1\) after you divide everything by \(\large t^2\)

OpenStudy (anonymous):

ohhhhh my goodness i see!

OpenStudy (anonymous):

huge error on my part. . . thanks so fricking much jim!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

and it wasn't so huge, silly typo really